Given the linear equation y=-5/9x-3 , find the y-coordinates of the points (-9, ), (0, ), and (9, ). Please show all of your work.
Answer by rfer(16322)

y=-5/9*-9-3
y=45/9-3
y=5-3
y=2
(-9,2)
------------------
y=-5/9*0-3
y=-3
(0,-3)
-------------
y=-5/9*9-3
y=-45/9-3
y=-5-3
y=-8
(9,-8)